Its about avogadro's law for gases..


If the ratio of masses in a container of O2 and N2 is 1:4. Then the ratio of there molecules will be..?



A) 7:32


B) 32:7


C) 4:1


D) 1:4


E) none

1
Expert's answer
2022-01-28T14:51:02-0500

Answer is (D) 1:4

Because, assuming you have 1g of O2 and 4g of N2,

Then moles will be

Moles of O2 =1/32 = 0.031mol

Moles of N2 = 4/28 = 0.142

Moles ratios = 0.031:0.142=1:4